.Suppose Ethiopian Electric Light and Power Corporation (EELPC) is a multi plantmonopolist having two plants, Tekeze plant (plant1) and Fincha plant (Plant2). Theoperating costs of the two plants are given as follows:Tekeze Plant: TC1 = 10 Q12and where Q1-Amount of electric power produced inTekezeFincha plant: TC2 = 20 Q22Q2–amount of electric power produced in FinchaEELPC estimates the demand for electric power by the following functionP= 700–5Q where P-is price (total in million birr) per Giga watt andQ–is the total amount of Giga watt sold and Q = Q1 + Q2


1
Expert's answer
2021-05-04T07:29:31-0400

"TC_1 = 10Q_1^{2 }"

"P_1 = 700-5Q _1"

"TR _1= P_1\u00d7Q_1 = (700-5Q_1)Q _1"

"TR_1 = 700Q_1 -5Q_1^{2 }"

Equilibrium point "MC = MR"

"MC_1 = 20Q_1"

"MR_1 = 700 - 10Q _1"

"20Q_1 = 700 - 10Q_1"

"30Q _1= 700"

"Q _1=23"

"P_1 = 700 - 5(Q_1)"

"P _1= 700 - 115"

"P_1=585"

"TC_1=10\u00d723^{2}=5290"

"MC_1=20\u00d723=460"

Therefore the equilibrium price is P1=585 and equilibrium quantity Q1 = 23


"TC_2 = 20Q_2^{2 }"

"P_2 = 700-5Q _2"

"TR _2= P_2\u00d7Q_2 = (700-5Q_2)Q _2"

"TR_2= 700Q_2 -5Q_2^{2 }"

Equilibrium point "MC = MR"

"MC_2= 40Q_2"

"MR_2= 700 - 10Q _2"

"40Q_2= 700 - 10Q_2"

"50Q _2= 700"

"Q _2=14"

"P_2 = 700 - 5(Q_2)"

"P _2= 700 - 70"

"P_2=630"

"TC_2=20\u00d714^{2}=3920"

"MC_2=40\u00d714=560"

Therefore the equilibrium price is P2=630 and equilibrium quantity Q 2= 14

total amount of Giga watt sold and

Q = Q1 + Q2

"Q=23+14=37"